Hello I need help with this question

During the budget period ended (31/12/2018), the total of revenues received was 530 million, rent paid was 200 million and Salaries paid was 190 million. In addition, the balance of treasury At the beginning of 2018 was 60 million. moreover, there are maintenance Contracts Which have been signed with service companies worth 10 millions JD and have not been implemented yet

Answer the following questions based on the Cash Basis method:


1. The balance of treasury at the end of 2018 is: 

A. 220

B. 200

C. 80

D. None of the above


2. The balance of the final account for the current period is: 

A. 100 Deficit

B. 40 Deficit

C. 140 Surplus

D. None of the above


3. The surplus at the end of 2018 is: 

A 220

B. 140

C. 20

D. None of the above


4. The balance of Commitment Reserve is:

A. 0

B. 200

C. 780

D. None of the above


1
Expert's answer
2022-02-03T12:16:26-0500

1.B

60+530-(200+190)=200


2.C

530-200-190=140


3.D

60+530-(200+190)=200


4.B

60+530-(200+190)=200